OpenStudy (anonymous):

anyone from CA with ALg 2. B

OpenStudy (anonymous):

simplify the trignometic expression: sin^2theta/1+costheta

OpenStudy (anonymous):

1.) sin(theta)+cos(theta)cot(theta) ---- sin(x) + cos(x*(cos/sin) ---- = (sin^2(x) + cos^2(x))/sin(x) = 1/sin(x) = csc(x) ======================================================= 2.)cos^2(theta)+tan^2(theta)cos^2(theta) ---- cos^2(x) + [sin^2(x)/cos^2(x)*cos^2(x)] ---- = cos^2(x) + sin^2(x) ----- = 1

OpenStudy (anonymous):

i dont have that in my answer choices: Ill show you a. cos theta b 1 + cos theta c. 1 - cos theta d. 1 - cos theta/ cos thetta

OpenStudy (anonymous):

ok ill try and solve that

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

im still checking the answer but it might be A but im not positive yet so dont put that

OpenStudy (anonymous):

ok where are you checing the answer

OpenStudy (anonymous):

what type of math is this exactly

OpenStudy (anonymous):

i mean like is it trigonometry